[Hallc_running] [New Logentry] RC Report -- 09/02/22

dhamilto at jlab.org dhamilto at jlab.org
Fri Sep 2 09:40:03 EDT 2022

Logentry Text:
We are currently in production running at 2.5 uA on LD2 for the final kinematic setting (Q2=1.60, W=3.08 Th_shms=6.00).

Heep-singles runs continued through the Day shift, with data-taking for all eight settings completed around 1pm. We then switched to single-arm HMS luminosity scans, which continued into the start of the Owl shift and were then followed by a series of SHMS lumi scans taken through the night. There were some CODA crashes and a bit of beam instability but no major issues. Thanks to a heroic effort by the respective shift crews (with more than 20 runs taken per shift), we were ready for the SHMS move to 6 degrees at 8am as planned. Steve Lassiter supervised remotely and has now disabled remote rotation of the spectrometers (https://logbooks.jlab.org/entry/4037012).

We will keep running at this setting for the remainder of the beamtime. 

Bob looked into an MSS file truncation problem and came up with a solution (https://logbooks.jlab.org/entry/4036503)
The dedicated BCM calibration using the Farday cup still needs to be scheduled. Doug suggested that immediately after the shutdown for maintenance/recovery would be a good time to do it.

Let's have a short meeting today at 11am.
Meeting ID: 160 710 0885 - Passcode: 595071


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4037040
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220902/efd7f659/attachment.html>

More information about the Hallc_running mailing list